If 2 is one of the eigenvalue of

−2 2 −3

2 1 −6

−1 −2 0

 then find the other two eigenvalues.


1
Expert's answer
2021-10-11T10:03:15-0400

"A=\begin{pmatrix}
 -2&2&-3 \\
 2&1&-6\\
-1&-2&0
\end{pmatrix}"

"|A-I\lambda|=\begin{vmatrix}
 -2-\lambda& 2& -3\\
 2& 1-\lambda & -6\\
 -1 & -2 & -\lambda
\end{vmatrix}"

"=-(\lambda-5)(\lambda+3)^2"

Equate to zero.

"\implies \lambda=5, -3"